OpenStudy (chasebraves):

3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

OpenStudy (wherearemytacos):

What are you confused about?

OpenStudy (chasebraves):

how to do it

OpenStudy (aihberkhan):

Okay this is very simply but very long, but I think I can help you out :) @chasebraves

OpenStudy (chasebraves):

thanks

OpenStudy (aihberkhan):

So first you want to distribute \(3(x-2)\), when you do distribute it you should get \(3x - 6\)

OpenStudy (aihberkhan):

So now your equation would look like \(3x - 6 + 4(2x - 6) = 6(x - 4) + 8(2x + 1)\)

OpenStudy (aihberkhan):

Next, we have to distribute \(4(2x-6)\), when you distribute you should get \(8x - 24\)

OpenStudy (aihberkhan):

So now your equation should look like \(3x - 6 + 8x - 24 = 6(x - 4) + 8(2x + 1)\)

OpenStudy (chasebraves):

ok

OpenStudy (aihberkhan):

So now that was just to give you an idea... now let me teach you ALL of it in a different away at once :)

OpenStudy (aihberkhan):

First you have to simplify \(3(x + -2) + 4(2x + -6) = 6(x + -4) + 8(2x + 1)\) Now lets try to reorder all of the terms \(3(-2 + x) + 4(2x + -6) = 6(x + -4) + 8(2x + 1)\) \((-2 * 3 + x * 3) + 4(2x + -6) = 6(x + -4) + 8(2x + 1)\) \((-6 + 3x) + 4(2x + -6) = 6(x + -4) + 8(2x + 1)\) We are not finished though. You want to keep reordering as much as you can \((-6 + 3x + 4(-6 + 2x) = 6(x + -4) + 8(2x + 1)\) \(-6 + 3x + (-6 * 4 + 2x * 4) = 6(x + -4) + 8(2x + 1)\) \(-6 + 3x + (-24 + 8x) = 6(x + -4) + 8(2x + 1)\) \(-6 + -24 + 3x + 8x = 6(x + -4) + 8(2x + 1)\) Now you want to combine the like terms \( -6 + -24 = -30\) That would become \(-30 + 3x + 8x = 6(x + -4) + 8(2x + 1)\) \(3x + 8x = 11x\) That would become \(-30 + 11x = 6(x + -4) + 8(2x + 1)\) Putting it all together: \(-30 + 11x = 6(-4 + x) + 8(2x + 1)\) \(-30 + 11x = (-4 * 6 + x * 6) + 8(2x + 1)\) \(-30 + 11x = (-24 + 6x) + 8(2x + 1)\) Now you keep following the steps until you can't simplify or reorder anymore. Then, you want to divide each side by \(-11\) This would give you \(x = -1.272727273\)

OpenStudy (aihberkhan):

So the final answer is \(x = -1.272727273\)

OpenStudy (aihberkhan):

Hope this helped! Have a great day! :) A medal would also be appreciated! Just click best response next to my name! Also, a fan would be appreciated as well! Just hover over my icon and click "Become A Fan". This will allow you to see every time I am online! :) If you see that I am online and need help with a question, just tag me in your question! @chasebraves

OpenStudy (chasebraves):

ok thanks